# Beyond Energy: Teleporting Current, Charge, and More

### v1

@@ Abstract
Quantum Energy Teleportation (QET) [5, 6, 7, 8], recently experimentally verified [9, 10, 11, 12], enables energy transfer between distant quantum parties using local operations and classical communication (LOCCs). Leveraging quantum entanglement and feedback control, QET is gaining attention in condensed matter, high energy physics, and quantum communication [13, 14, 15, 16, 17, 18, 19].
The aim of this work is to answer the questions below:
@@
    How is it related to entanglement?(Figs. 3,4)
@@
The CMD is fully consistent with information thermodynamics. Let $X$ be the system after the measurement and $X^{\prime}$ be the evolution of $X$ after feedback. The second law of the thermodynamics is equivalent to [35, 42]
@@
To provide an intuitive grasp of the inequality (8), we present the fidelity (7) in Fig. 2. The significant relationship with the teleported quantities is apparent in Fig. 1. Notably, the peaks in energy, electric current, and electric charge align with the minimal values of the fidelity. Moreover, this relationship underscores the global trend that lower fidelity results in more efficient teleportation, whereas higher fidelity leads to less teleportation, which gives another confirmation of eq. (8).
@@
Figure 3: Comparison between the ground state entanglement entropy $S_{EE}(\rho_{0})$ (left) and the charge susceptibility $\Delta\chi(Q)=\chi(Q,\rho_{\text{FC}})$ induced by the feedback-control (right).
Remarkably the diagram of $\Delta\chi(Q)$ captures the significant qualitative property of the phase diagram drawn by the ground state entanglement entropy $S_{EE}$ between the left and right subsystems. This agreement is extremely nontrivial because of the following reasons. First of all, $\Delta\chi(Q)$ is entirely determined by $\rho_{\text{FC}}$ since $\chi(Q,\rho_{0})=0$. However, as confirmed in Fig. 1, no teleportation occurs at $\mu_{5}a=1,m=0$, where the commutator of Bob’s operation and the Hamiltonian under the OBC vanishes:
$[U_{n_{B}},H_{\mu_{5}a=1,m=0}]=0.$ (13)
@@
To understand this better, we explore the current-current correlation function [48, 49]:
@@
In particular $\Delta\Pi^{00}_{ij}(t=0)$ is the spacial charge-charge correlation function and allows us to examine each term in the susceptibility (12). A result is plotted in Fig. 4 and more details including dynamics at $t>0$ are presented in SM and related to the operator growth [50, 51, 52]. The results would give us a new insight into the quantum correlation induced by LOCCs and quantum feedback-controls. The nontrivial aspects of the results are summarized below:
@@
: Asserting the upper bound of teleportation in its most general form (8).
@@
We emphasize that the findings presented in this work are anticipated to emerge across a diverse range of physical systems. A mid-circuit measurement of the ground state inherently elevates its energy, consequently inducing local currents and charges. Subsequently, control operations can instantaneously activate not only local energy but also currents and charges at distant locations. This process thereby instigates a novel flow of energy, current, charge, and other dynamic phenomena.
@@ I.1 Preliminaries
Figure 6: Outline of the setup, where Alice is the detector and Bob is the controller.
@@ I.2 Step 1: Deposit Energy
First, Alice measures her qubit in the ground state |$\ket{gs}$ of the system. Then the net amount she activates is evaluated by
@@
Since her measurement device physically acts on the system, it also injects the energy $\langle\Delta E_{0}\rangle$.
@@ II.1 Hamiltonian of the model
$H_{\text{obc}}=\frac{1}{4a}\sum_{n=1}^{N-1}(X_{n}X_{n+1}+Y_{n}Y_{n+1})+\frac{m}{2}\sum_{n=1}^{N}(-1)^{n}Z_{n}+\frac{\mu_{5}}{4}\sum_{n=1}^{N-1}(X_{n}X_{n+1}-Y_{n}Y_{n+1})$ (37)
$H_{\text{pbc}}=&\frac{1}{4a}\sum_{n=1}^{N-1}(X_{n}X_{n+1}+Y_{n}Y_{n+1})+\frac{m}{2}\sum_{n=1}^{N}(-1)^{n}Z_{n}+\frac{\mu_{5}}{4}\sum_{n=1}^{N-1}(X_{n}X_{n+1}-Y_{n}Y_{n+1})\\ &+\frac{(-1)^{\frac{N}{2}}}{4a}\left[\left(X_{N}X_{1}+Y_{N}Y_{1}\right)+\mu_{5}\left(X_{N}Y_{1}-Y_{N}X_{1}\right)\right]\prod_{n=2}^{N-1}Z_{n}.\end{aligned}$ (38)
@@ II.3 Charge susceptibility
Figure 10: Left: the entanglement entropy. Right: the charge susceptibility. $N=8$ under the PBC.
@@ II.4 Real-time evolution of the charge-charge correlation function
Figure 12: Real-time evolution of the current-current correlation function $\Delta\Pi^{00}_{0n_{B}}(t)$. The simulation was performed with parameters $ma=0.1,\mu=5a=1,N=16$ under the OBC. Similar plots can be obtained with the PBC.
@@ III.2 Quantum Current Teleportation
The massive Schwinger model exhibits a critical point at $m/g\approx 0.33$, around which the teleported current displays an intriguing behavior, as illustrated in Fig. 13. Our simulation was conducted with an $N=8$ lattice under the OBC.

### v2

@@ Abstract
Quantum Energy Teleportation (QET) [5, 6, 7, 8], recently experimentally verified [9, 10, 11, 12], enables energy transfer between distant quantum parties using local operations and classical communications (LOCCs). Leveraging quantum entanglement and feedback control, QET is gaining attention in condensed matter, high energy physics, and quantum communication [13, 14, 15, 16, 17, 18, 19].
We aim to answer the following open questions:
@@
    How is it related to entanglement?(Figs. 3,4&SM)
@@
The CMD is fully consistent with information thermodynamics. Let $X$ be the system after the measurement and $X^{\prime}$ be the evolution of $X$ after feedback. The second law of thermodynamics is equivalent to [35, 42]
@@
To provide an intuitive grasp of the inequality (8), we present the fidelity (7) in Fig. 2. The significant relationship with the teleported quantities is apparent in Fig. 1. Notably, the peaks in energy, electric current, and electric charge align with the minimal values of the fidelity. Moreover, this relationship underscores the global trend that lower fidelity results in more efficient teleportation, whereas higher fidelity leads to less teleportation, which gives another confirmation of eq. (8). See also SM.
@@
Figure 3: Comparison between the ground state entanglement entropy $S_{A\bar{A}}(\rho_{0})$ (left) and the charge susceptibility $\Delta\chi(Q)=\chi(Q,\rho_{\text{FC}})$ induced by the feedback-control (right).
Remarkably the diagram of $\Delta\chi(Q)$ captures the significant qualitative property of the phase diagram drawn by the ground state entanglement entropy $S_{A\bar{A}}(\rho_{0})$ between Alice $A$ and its complement $\bar{A}$. This agreement is extremely nontrivial because of the following reasons. First of all, $\Delta\chi(Q)$ is entirely determined by $\rho_{\text{FC}}$ since $\chi(Q,\rho_{0})=0$. However, as confirmed in Fig. 1, teleportation is apparently not directly related to entanglement, although its upper bound is related to the entanglement entropy and fidelity, as discussed in SM. In fact, $\Delta\chi(Q)$ holds a large value even at a point where teleportation does not occur at all.
@@
To understand this phenomenon better, we explore the current-current correlation function [48, 49]:
@@
In particular $\Delta\Pi^{00}_{ij}(t=0)$ is the spatial charge-charge correlation function and allows us to examine each term in the susceptibility (12). A result is plotted in Fig. 4 and more details including dynamics at $t>0$ are presented in SM and related to the operator growth [50, 51, 52]. The results would give us a new insight into the quantum correlation induced by LOCCs and quantum feedback-controls. The nontrivial aspects of the results are summarized below:
@@
: Asserting the most general upper bound of teleportation (8). Generally this could be a stronger upper bound than the known bound given by the entanglement entropy $S_{A\bar{A}}(\rho_{0})$ [69], which can exceed 1. The formula is further generalized in SM.
@@
We emphasize that our findings are expected to appear in various physical systems. A mid-circuit measurement of the ground state raises its energy, inducing local currents and charges. Control operations can then instantaneously activate energy, currents, and charges at distant locations, triggering novel dynamic phenomena.
@@ I.1 Preliminaries
Figure 6: Outline of the setup, where Alice at $n_{A}$ is the detector and Bob at $n_{B}$ is the controller. They are spatially separated and communicate through classical channel.
@@ I.2 Step 1: Deposit Energy
First, Alice measures her qubit in the ground state |$\ket{gs}$ of the system. Then the net average amount she activates is evaluated by
@@
Since her measurement device physically acts on the system, it also injects the energy $\langle\Delta E_{0}\rangle$ into the system.
@@ I.4 Comment on the Maxwell Demon and QET
I.5 Upper bound on teleportation: Entropy vs Fidelity
Here we give a general remark on the upper bound of teleportation (8). So far, the known best upper bound was given by the entanglement entropy between Alice’s subsystem and its complement $S_{A\bar{A}}(\rho_{0})$ [69] in such a way that
$\frac{E^{2}_{n_{B}}}{4\|H_{n_{B}}\|^{2}}\leq S_{A\bar{A}}(\rho).$ (33)
However, in general, the upper bound of $S_{A\bar{A}}(\rho_{0})$ is determined by the area of $A$ (the number of qubits in $A$), the dimension of the system and the energy gap between the lowest energy and the 1st excited energy. Therefore $S_{A\bar{A}}$ can be larger than 1, depending on the systems and the parameters. In particular, when a system undergoes a quantum phase transition, it diverges at a critical point and becomes very large around it. Of course, the energy is finite around such a critical point, so the teleported energy is much lower than the entanglement entropy. In fact, the theoretically rigorous upper bound given by the fidelity $1-F(\rho_{0},\rho_{\text{FC}})$ (8) guarantees that the ratio $E_{n_{B}}/4\|H_{n_{B}}\|$ does not exceed 1. A closer look at Figs. 2 and 3 clarifies that $1-F(\rho_{0},\rho_{\text{FC}})$ can be greater or smaller than $S_{A\bar{A}}(\rho_{0})$, depending on the points.
Here we obtain the state-of-the-art upper bound of the teleported quantity. For a general observable $\mathcal{O}_{n_{B}}$ at Bob’s site, we can still apply the discussion in [69]:
$\frac{|\langle\Delta\mathcal{O}_{n_{B}}\rangle-\langle\mathcal{O}_{n_{B}}\rangle|^{2}}{4\|\mathcal{O}_{n_{B}}\|^{2}}\leq S_{A\bar{A}}(\rho_{0}),$ (34)
where $\langle\mathcal{O}_{n_{B}}\rangle=\sum_{b}p_{A}(b)\mathrm{Tr}[\rho_{0}U^{\dagger}_{n_{B}}(b)\mathcal{O}U_{n_{B}}(b)]$ with the probability $p_{A}(b)$ that $b$ is observed by Alice. When the sign of $\langle\Delta\mathcal{O}_{n_{B}}\rangle$ is opposite to the sign of $\langle\mathcal{O}_{n_{B}}\rangle$, we obtain a new upper bound:
$\frac{|\langle\Delta\mathcal{O}_{n_{B}}\rangle|^{2}}{4\|\mathcal{O}_{n_{B}}\|^{2}}\leq\min\{S_{A\bar{A}}(\rho_{0}),1-F(\rho_{0},\rho_{\text{FC}})\}.$ (35)
In particular, when $O_{n_{B}}$ is Bob’s local Hamiltonian, the previous upper bound (33) is now updated as follows:
$\frac{E^{2}_{n_{B}}}{4\|H_{n_{B}}\|^{2}}\leq\min\{S_{A\bar{A}}(\rho_{0}),1-F(\rho_{0},\rho_{\text{FC}})\}.$ (36)
@@ II.1 Hamiltonian of the model
$H_{\text{obc}}=\frac{1}{4a}\sum_{n=1}^{N-1}(X_{n}X_{n+1}+Y_{n}Y_{n+1})+\frac{m}{2}\sum_{n=1}^{N}(-1)^{n}Z_{n}+\frac{\mu_{5}}{4}\sum_{n=1}^{N-1}(X_{n}Y_{n+1}-Y_{n}X_{n+1})$ (40)
$H_{\text{pbc}}=&\frac{1}{4a}\sum_{n=1}^{N-1}(X_{n}X_{n+1}+Y_{n}Y_{n+1})+\frac{m}{2}\sum_{n=1}^{N}(-1)^{n}Z_{n}+\frac{\mu_{5}}{4}\sum_{n=1}^{N-1}(X_{n}Y_{n+1}-Y_{n}X_{n+1})\\ &+\frac{(-1)^{\frac{N}{2}}}{4a}\left[\left(X_{N}X_{1}+Y_{N}Y_{1}\right)+\mu_{5}\left(X_{N}Y_{1}-Y_{N}X_{1}\right)\right]\prod_{n=2}^{N-1}Z_{n}.\end{aligned}$ (41)
@@ II.3 Charge susceptibility
Figure 10: Left: the ground state entanglement entropy between Alice’s system $A$ and its complement $\bar{A}$. Right: the charge susceptibility. $N=8$ under the PBC.
@@ II.4 Real-time evolution of the charge-charge correlation function
Figure 12: Real-time evolution of the charge-charge correlation function $\Delta\Pi^{00}_{0n_{B}}(t)$. The simulation was performed with parameters $ma=0.1,\mu=5a=1,N=16$ under the OBC. Similar plots can be obtained with the PBC.
@@ III.2 Quantum Current Teleportation
The massive Schwinger model exhibits a critical point at $m/g\approx 0.33$, around which the teleported current displays an intriguing behavior, as illustrated in Fig. 13. Our simulation was conducted with an $N=8$ lattice under the OBC. Here we used $\phi$ that is given by eq. (24), although the model is all-to-all interacting and thus $[H_{B},\sigma_{B}]\neq[H,\sigma_{B}]$.